four fifths of a number x when added to two fifths of the same no. x gives result 100. Find x?
Answer by stanbon(75887)

four fifths of a number x when added to two fifths of the same no. x gives result 100. Find x?
----
(4/5)x + (2/5)x = 100
6/5 x = 100
x = (5/6)100
x = 500/6 = 83 1/3
-----------------
Cheers,
Stan H.
--------------